Krel:

The sides of a triangle are in the ratio of 1 : 2 : 2 and the perimeter of the triangle is 15 feet. What´s the length of the shortest side?

Mathmachine2021:

hi

Krel:

Hi.

Mathmachine2021:

i think 3 feet

Mathmachine2021:

do i get a meadl now

Krel:

Thank you.

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mathmachine2021 do i get a meadl now \(\color{#0cbb34}{\text{End of Quote}}\) no you need to help them figure it out not guess

xxDeppressionxx:

1x + 2x + 2x = 15 5x = 15 x = 15/5 x = 3

Mathmachine2021:

ohhhhhhhhh

Mathmachine2021:

im new so i didnt know sorry

Mathmachine2021:

:(

Krel:

Thanks!

Krel:

What metal?